You cross two true-breeding lines of flies, one line with normal eyes and the other with sepia eyes. All F1 flies produced have normal eyes. If the F1’s are crossed to other F1’s and a total of 1,810 baby flies are produced, how many do you expect to have normal eyes if this trait is due to a single gene showing complete dominance?
Answer #1

Let S be the dominant allele for normal eyes

Let s be the recessive allele for sepia eyes

Parent cross: SS (normal eyes) * ss (sepia eyes)

Gametes of normal eyed parent: S

Gametes of sepia eyed parent: s

F1 offspring: Ss (normal eyes)

F1 cross: Ss (norma eyes) * Ss (normal eyes)

Gametes of F1 male: S, s

Gametes of F1 female: S, s

F2 offspring: SS (normal), Ss (normal), Ss (normal), ss (sepia)

Genotypic ratio: 1(SS) : 2(Ss) : 1(ss)

Phenotypic ratio: 3(normal eyes) :1(sepia eyes)

Probability of F2 offspring with normal eyes = 3/4 or 75%

Probability of F2 offspring with sepia eyes = 1/4 or 25%

Total offspring = 1810

Expected number of offspring with normal eyes = probability % * total offspring = 0.75 * 1810 = 1357.5

Therefore, expected number of offspring with normal eyes = 1357.5 = 1358 (rounded)